# The DESI 2024 hint for dynamical dark energy is biased by low-redshift supernovae

### v3

@@ Abstract
Recently, a $\sim 3.9\sigma$ preference for dynamical dark energy has been reported by the Dark Energy Spectroscopic Instrument (DESI) collaboration, which has inspired hot debates on new physics or systematics. In this letter, we reveal that this significant preference is dominated by an external low-redshift supernova (low-$z$ SN) sample, which was combined with the Dark Energy Survey SN program (DES-SN) in their Year-Five data release (DESY5). Using the intercept in the SN magnitude-redshift relation as a diagnostics for systematics, we find not only large dispersions but also a large discrepancy in the low-$z$ SN sample when compared to the high-$z$ DES-SN sample within the single DESY5 compilation. Correcting for this low-$z$ systematics with or without including the cosmic microwave background data can largely reduce the preference for dynamical DE to be $0.5\sim 1.5\sigma$. Therefore, the DESI preference for dynamical DE might be a mirage of unknown systematics in the low-$z$ SN sample.
@@ 1 Introduction
As one of the stage-III surveys, the Dark Energy Survey (DES) explores the properties of dark matter and DE with unprecedented precision and accuracy through multiple probes like Type Ia supernovae (SNe Ia), weak lensings, large-scale structures, galaxy clusters (T. Abbott et al., 2005; J. P. Bernstein et al., 2012; T. Abbott et al., 2016). As the most updated Stage IV survey, the recent release of the Dark Energy Spectroscopic Instrument (DESI) for baryon acoustic oscillations (BAO) (A. G. Adame et al., 2024a) together with full 5-year photometrically classified 1635 SNe Ia (DES-SN), after combined with previous external 194 low-$z$ SNe Ia (low-$z$ SN) in the DES Year-Five Data Release (DESY5) (T. M. C. Abbott et al., 2024; M. Vincenzi et al., 2024), have triggered heated debates on the existence of dynamical DE. Specifically, the joint constraint of cosmic microwave background (CMB) Planck data, DESI BAO measurements, and DESY5 (low-$z$ SN+DES-SN) compilation surprisingly favors a quintom-like dynamical DE behavior, showing a $3.9\sigma$ tension with the $\Lambda$CDM model (A. G. Adame et al., 2024a, b). A similar but weaker preference ($\sim 2\sigma$: C.-G. Park et al., 2024a) of dynamic DE is also achieved without the participation of DESI BAO and DESY5. These hints for new physics have attracted significant interests in the community with subsequent numerous theoretical studies on different kinds of dynamical DE models (e.g., Y. Tada & T. Terada, 2024; K. V. Berghaus et al., 2024; D. Shlivko & P. J. Steinhardt, 2024; O. F. Ramadan et al., 2024; S. Bhattacharya et al., 2024; W. Giarè et al., 2024; I. J. Allali et al., 2024; K. S. Croker et al., 2024; Y. Yang et al., 2024; C. Escamilla-Rivera & R. Sandoval-Orozco, 2024; W. Yin, 2024; G. Gu et al., 2024; X. Wang et al., 2024; M. Cortês & A. R. Liddle, 2024; E. O. Colgáin et al., 2024a; Y. Carloni et al., 2024; K. Lodha et al., 2024; R. Calderon et al., 2024; O. Luongo & M. Muccino, 2024; P. Mukherjee & A. A. Sen, 2024; B. R. Dinda, 2024; R. Camilleri et al., 2024; T.-N. Li et al., 2024; Q. Gao et al., 2024; R. Fikri et al., 2024; T.-N. Li et al., 2025).
@@
In parallel with theoretical explanations, inconsistencies between recent new measurements (DESI BAO, DESY5) and previous high-quality measurements, such as SDSS BAO (S. Alam et al., 2017, 2021) and PantheonPlus SNe (A. G. Riess et al., 2022; D. Brout et al., 2022a; D. Scolnic et al., 2022; D. Brout et al., 2022b; E. R. Peterson et al., 2021; A. Carr et al., 2022; B. Popovic et al., 2023), have also been widely debated (e.g., C.-G. Park et al., 2024b; C.-G. Park & B. Ratra, 2025; G. Liu et al., 2024; E. O. Colgáin et al., 2024b; E. O. Colgáin & M. M. Sheikh-Jabbari, 2024; Z. Wang et al., 2024; Z. Huang et al., 2024; S. Dhawan et al., 2024; S. Roy Choudhury & T. Okumura, 2024; I. D. Gialamas et al., 2024; G. Efstathiou, 2024; A. Notari et al., 2024; Z. Sakr, 2025a, b). For example, the combination of DESI LRG1 and LRG2 data is found to be the key driver for the preference of dynamical DE (G. Liu et al., 2024; Z. Wang et al., 2024; Z. Huang et al., 2024; A. G. Adame et al., 2024a; E. O. Colgáin et al., 2024a). Nevertheless, these two data exhibit mild tension with SDSS BAO (S. Alam et al., 2021) for unknown reasons. What’s more, by comparing the mean magnitude-redshift residuals between PantheonPlus and DESY5, G. Efstathiou (2024) found a $\sim 0.04-0.05$ mag. difference in the overlapping regime, indicating unidentified SNe systematics in DESY5 (see, however, M. Vincenzi et al., 2025). Correcting this offset would reduce the preference for dynamical DE, bringing cosmological constraints of the corrected DESY5 back to $\Lambda$CDM. Therefore, it cannot simply rule out the possibility of potential systematics in DESY5.
@@
In this letter, we first cross-compare DESI BAO and DESY5 SNe with the previous high-quality SDSS eBOSS BAO and PantheonPlus SNe, and confirm that the key driver of dynamical DE is the additional external low-$z$ SN sample in forming the DESY5 compilation. We then propose to diagnose the systematics of DESY5 and quantify its preference for dynamic DE through the dubbed $a_{B}$ consistency developed by L. Huang et al. (2025, 2024). By performing simple corrections to DESY5, we find the fully de-biased DESY5 SNe, after combining with DESI BAO regardless of CMB, shows no significant deviation from the $\Lambda$CDM model. Therefore, it is too early to claim dynamical DE.
Figure 1: Left: The $w_{0}-w_{a}$ constraits from CMB+BAO(eBOSS/DESI)+SNe(PantheonPlus/DESY5). Right: The $w_{0}-w_{a}$ constraints from CMB+DESY5/DES-SN/low-$z$ (with and without fluctuations debiased).
@@ 2 Data and method
As indicated by A. G. Adame et al. (2024a, b), different late-time observations seem to have different levels of support for the dynamical DE. To diagnose which dataset contributes the most to the preference for dynamical DE, we take all the high-quality DESI BAO (A. G. Adame et al., 2024a), DESY5 (low-$z$ SN + DES-SN: T. M. C. Abbott et al., 2024; M. Vincenzi et al., 2024), SDSS eBOSS BAO (S. Alam et al., 2021), PantheonPlus (A. G. Riess et al., 2022; D. Brout et al., 2022a; D. Scolnic et al., 2022; D. Brout et al., 2022b; E. R. Peterson et al., 2021; A. Carr et al., 2022; B. Popovic et al., 2023), and Planck CMB (N. Aghanim et al., 2020) into considerations for completed comparisons. The UNION3 compilation (D. Rubin et al., 2023) is not included in this study as they have not released a Hubble diagram yet.
@@
Similar to PantheonPlus SNe analysis, we construct the distance moduli residual vector of DESY5 SNe as $\Delta\vec{D}=m_{B}^{\mathrm{corr.}}-M_{B}-\mu_{\mathrm{model}}$, where standardized apparent magnitudes $m_{B}^{\mathrm{corr.}}$ have already corrected for stretch, color, simulation bias (selection bias, Malmquist bias, and light curve fitting bias), and mass step effects. The likelihood is defined as $-2\ln(\mathcal{L})=\Delta\vec{D^{T}}C^{-1}_{\mathrm{stat+syst}}\Delta\vec{D}$. We adopt the commonly used builtin Planck CMB (plik TTTEEE + lowl + lowE + CMB lensing), SDSS eBOSS BAO (eight BAO-only data points), and PantheonPlus (local SNe at $z<0.01$ are discarded) in MontePython (B. Audren et al., 2013), along with extra self-written DESI BAO and DESY5 likelihoods. In what follows, we perform the Monte Carlo Markov Chain (MCMC) on the $w_{0}w_{a}$ model with Planck CMB + arbitrarily different combinations of BAO and SNe to locate the possible systematics direction.

### v4

# The DESI DR1/DR2 evidence for dynamical dark energy is biased by low-redshift supernovae

@@ Abstract
Recently, a $3\sim 4\sigma$ preference for dynamical dark energy has been reported by the Dark Energy Spectroscopic Instrument (DESI) collaboration, which has inspired hot debates on new physics or systematics. In this letter, we reveal that this preference is significantly biased by an external low-redshift supernova (low-$z$ SN) sample, which was combined with the Dark Energy Survey SN program (DES-SN) in their Year-Five data release (DESY5). Using the intercept in the SN magnitude-distance relation as a diagnostic for systematics, we find not only large dispersions but also a large discrepancy in the low-$z$ SN sample when compared to the high-$z$ DES-SN sample within the single DESY5 compilation. Correcting for this low-$z$ systematics with or without including the cosmic microwave background data can largely reduce the preference for dynamical DE to be less than $2\sigma$. Therefore, the DESI preference for dynamical DE is biased by some unknown systematics in the low-$z$ SN sample.
@@ 1 Introduction
As one of the stage-III surveys, the Dark Energy Survey (DES) explores the properties of dark matter and DE with unprecedented precision and accuracy through multiple probes like Type Ia supernovae (SNe Ia), weak lensings, large-scale structures, galaxy clusters (T. Abbott et al., 2005; J. P. Bernstein et al., 2012; T. Abbott et al., 2016). As the most updated Stage IV survey, the recent release of the Dark Energy Spectroscopic Instrument (DESI) for baryon acoustic oscillations (BAO) (A. G. Adame et al., 2025a, b, c) together with full 5-year photometrically classified 1635 SNe Ia (DES-SN), after combined with previous external 194 low-$z$ SNe Ia (low-$z$ SN) in the DES Year-Five Data Release (DESY5) (T. M. C. Abbott et al., 2024; M. Vincenzi et al., 2024), have triggered heated debates on the existence of dynamical DE. Specifically, the joint constraint of cosmic microwave background (CMB) Planck data, DESI DR1 (Y1 BAO measurements), and DESY5 (low-$z$ SN+DES-SN) compilation surprisingly favors a dynamical DE with phantom crossing, showing a $3.9\sigma$ tension with the $\Lambda$CDM model (A. G. Adame et al., 2025a, 2024). A similar but weaker preference ($\sim 2\sigma$: C.-G. Park et al., 2024a) of dynamic DE is also achieved without the participation of DESI DR1 and DESY5. These hints for new physics have attracted significant interests in the community with subsequent numerous theoretical studies on different kinds of dynamical DE models (e.g., Y. Tada & T. Terada, 2024; K. V. Berghaus et al., 2024; D. Shlivko & P. J. Steinhardt, 2024; O. F. Ramadan et al., 2024; S. Bhattacharya et al., 2024; W. Giarè et al., 2024; I. J. Allali et al., 2024; K. S. Croker et al., 2024; Y. Yang et al., 2024; C. Escamilla-Rivera & R. Sandoval-Orozco, 2024; W. Yin, 2024; G. Gu et al., 2024; X. Wang et al., 2024; M. Cortês & A. R. Liddle, 2024; E. O. Colgáin et al., 2024a; Y. Carloni et al., 2024; K. Lodha et al., 2024; R. Calderon et al., 2024; O. Luongo & M. Muccino, 2024; P. Mukherjee & A. A. Sen, 2024; B. R. Dinda, 2024; R. Camilleri et al., 2024; T.-N. Li et al., 2024; Q. Gao et al., 2024; R. Fikri et al., 2024; T.-N. Li et al., 2025).
@@
In parallel with theoretical explanations, inconsistencies between recent new measurements (DESI DR1, DESY5) and previous high-quality measurements, such as SDSS BAO (S. Alam et al., 2017, 2021) and PantheonPlus SNe (A. G. Riess et al., 2022; D. Brout et al., 2022a; D. Scolnic et al., 2022; D. Brout et al., 2022b; E. R. Peterson et al., 2021; A. Carr et al., 2022; B. Popovic et al., 2023), have also been widely debated (e.g., C.-G. Park et al., 2024b; C.-G. Park & B. Ratra, 2025; G. Liu et al., 2024; E. O. Colgáin et al., 2024b; E. O. Colgáin & M. M. Sheikh-Jabbari, 2024; Z. Wang et al., 2024; Z. Huang et al., 2024; S. Dhawan et al., 2024; S. Roy Choudhury & T. Okumura, 2024; I. D. Gialamas et al., 2024; G. Efstathiou, 2024; A. Notari et al., 2024; Z. Sakr, 2025a, b). For example, the combination of DESI LRG1 and LRG2 data is found to be the key driver for the preference of dynamical DE (G. Liu et al., 2024; Z. Wang et al., 2024; Z. Huang et al., 2024; A. G. Adame et al., 2025a; E. O. Colgáin et al., 2024a). Nevertheless, these two data exhibit mild tension with SDSS BAO (S. Alam et al., 2021) for unknown reasons. What’s more, by comparing the mean magnitude-distance residuals between PantheonPlus and DESY5, G. Efstathiou (2024) found a $\sim 0.04-0.05$ magnitude (mag.) difference in the overlapping regime, indicating unidentified SNe systematics in DESY5 (see, however, M. Vincenzi et al., 2025). Correcting this offset would reduce the preference for dynamical DE, bringing cosmological constraints of the corrected DESY5 back to $\Lambda$CDM. Therefore, it cannot simply rule out the possibility of potential systematics in DESY5.
@@
In this letter, we first cross-compare DESI DR1 and DESY5 SNe with the previous high-quality SDSS eBOSS BAO and PantheonPlus SNe, and confirm that the key driver of dynamical DE is the additional external low-$z$ SN sample in forming the DESY5 compilation. We then propose to diagnose the systematics of DESY5 and quantify its preference for dynamic DE through the dubbed $a_{B}$ consistency developed by L. Huang et al. (2025, 2024). By performing simple corrections to DESY5, we find the fully de-biased DESY5, after combining with DESI DR1/DR2 regardless of CMB, shows no significant deviation from the $\Lambda$CDM model. Therefore, it is too early to claim dynamical DE.
@@ 2 Data and method
Figure 1: Left: The $w_{0}-w_{a}$ constraits from CMB (Planck 2018) + BAO (eBOSS DR16/DESI DR1) + SNe (PantheonPlus/DESY5). Right: The $w_{0}-w_{a}$ constraints from CMB+DESY5/DES-SN/low-$z$ (with and without fluctuations debiased).
@@
As indicated by A. G. Adame et al. (2025a, 2024), different late-time observations seem to have different levels of support for the dynamical DE. To diagnose which dataset contributes the most to the preference for dynamical DE, we take all the high-quality DESI DR1 (A. G. Adame et al., 2025a), DESY5 (low-$z$ SN + DES-SN: T. M. C. Abbott et al., 2024; M. Vincenzi et al., 2024), SDSS eBOSS BAO (S. Alam et al., 2021), PantheonPlus (A. G. Riess et al., 2022; D. Brout et al., 2022a; D. Scolnic et al., 2022; D. Brout et al., 2022b; E. R. Peterson et al., 2021; A. Carr et al., 2022; B. Popovic et al., 2023), and Planck CMB (N. Aghanim et al., 2020) into considerations for completed comparisons. The UNION3 compilation (D. Rubin et al., 2023) is not included in this study as they have not released a Hubble diagram yet.
@@
Similar to PantheonPlus SNe analysis, we construct the distance moduli residual vector of DESY5 SNe as $\Delta\vec{D}=m_{B}^{\mathrm{std}}-M_{B}-\mu_{\mathrm{model}}$, where standardized apparent magnitudes $m_{B}^{\mathrm{std}}$ have already corrected for stretch, color, simulation bias (selection bias, Malmquist bias, and light curve fitting bias), and mass step effects. The likelihood is defined as $-2\ln(\mathcal{L})=\Delta\vec{D^{T}}C^{-1}_{\mathrm{stat+syst}}\Delta\vec{D}$. We adopt the commonly used built-in Planck CMB (plik TTTEEE + lowl + lowE + CMB lensing), SDSS eBOSS BAO (eight BAO-only data points), and PantheonPlus (local SNe at $z<0.01$ are discarded) in MontePython (B. Audren et al., 2013), along with extra self-written DESI BAO and DESY5 likelihoods. In what follows, we perform the Monte Carlo Markov Chain (MCMC) on the $w_{0}w_{a}$ model with Planck CMB + arbitrarily different combinations of BAO and SNe to locate the possible systematics direction.
3 Analysis
3.1 Systematics identification
In the Appendix, we present all constraints of the $w_{0}w_{a}$ model of this study in Tab. 1 and Tab. 2. The Tab. 1 is visualized in the left panel Fig. 1 for their marginalized $w_{0}-w_{a}$ distributions, where the base preference for dynamical DE from Planck-CMB+eBOSS-BAO+PantheonPlus is $\sim 1.7\sigma$ (blue contour in the upper left), not significant yet. Either replacing the eBOSS BAO with DESI DR1 (green contour in the upper right) or the PantheonPlus SNe with DESY5 SNe (purple contour in the bottom left) would increase preferences to $\sim 2.5\sigma$ and $\sim 2.9\sigma$, respectively, but still not decisive to claim dynamical DE. In particular, when getting rid of the low-$z$ SN sample, the significance drops from $\sim 2.9\sigma$ to insignificant $\sim 1.1\sigma$ (orange contour in the bottom left). Even though CMB+DESI DR1+DESY5 after both of the above replacements presents a more significant $\sim 3.5\sigma$ preference (red contour in the bottom right), getting rid of the low-$z$ SN sample would once again drop the significance from $\sim 3.5\sigma$ back to $\sim 1.7\sigma$ (orange dashed contour in the bottom right). On the contrary, both cases using low-$z$ SNe (blue dot-dashed contours in the bottom row) prefer a deviation from the $\Lambda$CDM model with a direction towards $w_{0}>-1$ and $w_{a}<0$. Therefore, a minor inconsistency emerges between low-$z$ SNe and DES-SN samples.
3.2 Systematics diagnosis
Figure 3: Left: The marginalized contours on parameters $w_{0}$ and $w_{a}$ for Planck CMB+DESI BAO+DESY5 with different de-biasing corrections. Right: the same as the left panel but excluding Planck-CMB with a fiducial $H_{0}=67.4$ km/s/Mpc. The DESI DR1 (Y1 BAO) and DR2 (Y3 BAO) are used in the top and bottom rows, respectively.
To diagnose systematics for low-$z$ SNe, we rebuild the magnitude-(logarithmic) distance relation
$m_{B,i}^{\mathrm{std}}$ $=5\lg\hat{d}_{L}(z_{i})-5a_{B},$ (1)
$-5a_{B}$ $\equiv M_{B}+5\lg\frac{c/H_{0}}{\mathrm{Mpc}}+25$ (2)
for both low-$z$ and DES-SN samples, where the dimensionless luminosity distance $\hat{d}_{L}(z_{i})=D_{L}(z)/(c/H_{0})=(1+z_{\mathrm{hel}})\int_{0}^{z_{\mathrm{HD}}}\mathrm{d}z^{\prime}/E(z^{\prime})$ is computed from the best-fit Planck $\Lambda$CDM fiducial model for the reduced Hubble parameter $E(z)^{2}\equiv H(z)^{2}/H_{0}^{2}=\Omega_{m}(1+z)^{3}+(1-\Omega_{m})$. Recall that all light-curve model corrections are used to standardize the apparent magnitude $m_{B,i}^{\mathrm{std}}$, and hence the intercept $-5a_{B}$ as a degeneracy parameter between $H_{0}$ and $M_{B}$ is expected to be a stable value for a single SNe compilation. The gist of $a_{B}$ diagnosis is to check whether the intercept $-5a_{B}$ (a systematics-sensitive and model-sensitive quantity) is consistent among different SNe samples. As any observational systematics in measuring $m_{B,i}^{\mathrm{std}}$ or theoretical systematics in modeling $5\lg\hat{d}_{L}(z_{i})$ induced by either uncounted peculiar velocity effect on the redshift $z_{i}$ measurement or even new physics model in $\hat{d}_{L}(z)$ would cause a deviation in the intercept $-5a_{B,i}=m_{B,i}^{\mathrm{std}}-5\lg\hat{d}_{L}(z_{i})$, we can quickly locate the redshift ranges of either systematics or new physics from redshift distributions of $-5a_{B}$.
Detailed $a_{B}$ diagnosis is visualized in Fig. 2. The high-quality DES-SN from a single telescope has the most stable $a_{B}$ distributions among its spanned redshift bins, perfectly agreeing with the Planck $\Lambda$CDM cosmology assumed in Fig. 2. By contrast, the multi-sample mixed low-$z$ SNe shows irregular fluctuations in $-5a_{B}$, and its weighted average is $\sim 0.043$ mag. higher than that of DES-SN, which is qualitatively similar to the previous study by G. Efstathiou (2024) but with the $\sim 0.04$ mag. offset estimated on $\Delta m_{\mathrm{offset}}:=(m_{\mathrm{Pantheon+}}^{\mathrm{low}-z}-m_{\mathrm{DESY5}}^{\mathrm{low}-z})-(m_{\mathrm{Pantheon+}}^{\mathrm{high}-z}-m_{\mathrm{DESY5}}^{\mathrm{high}-z})\approx(-0.05)-(-0.01)=-0.04$ mag. between PantheonPlus and DESY5 compilations in their low- and high-redshift ranges after assuming a constant $M_{B}$. As a response, M. Vincenzi et al. (2025) has explained partially this offset with improvements to the intrinsic scatter model and mass step estimate by reducing the offset from $\sim 0.04$ mag. to $\Delta\mu_{\mathrm{offset}}:=(\mu_{\mathrm{Pantheon+}}^{\mathrm{low}-z}-\mu_{\mathrm{DESY5}}^{\mathrm{low}-z})-(\mu_{\mathrm{Pantheon+}}^{\mathrm{high}-z}-\mu_{\mathrm{DESY5}}^{\mathrm{high}-z})=(\mu_{\mathrm{Pantheon+}}^{\mathrm{low}-z}-\mu_{\mathrm{Pantheon+}}^{\mathrm{high}-z})-(\mu_{\mathrm{DESY5}}^{\mathrm{low}-z}-\mu_{\mathrm{DESY5}}^{\mathrm{high}-z})=0.019-0.043=-0.024$ mag., which can be further reduced to $-0.008$ mag. after considering a $\sim 0.016$ mag. difference in selection functions as elaborated in Table 1 of M. Vincenzi et al. (2025). However, such a negligible $-0.008$ mag. offset in $\Delta\mu_{\mathrm{offset}}$ is actually achieved between two different SN compilations, leaving an equivalent $0.019+0.008\sim 0.027$ mag. difference between low-$z$ and high-$z$ SNe within the single DESY5 compilation, which is still apparent in the difference of $-5a_{B}$ between low-$z$ and high-$z$ parts of the single DESY5 compilation. Therefore, the key difference of our current study from the previous two analyses (G. Efstathiou, 2024; M. Vincenzi et al., 2025) is that they considered the low/high-redshift difference among two different SN compilations (PantheonPlus and DESY5) while we focus on the low/high-redshift difference within the same SN compilation.
This $\sim 0.043$ mag. discrepancy in $-5a_{B}$ within the single DESY5 compilation appears to emerge suddenly around $z\sim 0.1$ where the Universe is still homogeneous, hence we can safely rule out the local inhomogeneous effects. This sudden transition also cannot originate from homogeneous-scale new physics since the PantheonPlus compilation, containing a larger number of low-$z$ well-calibrated SNe, does not admit any such low-$z$ transition. If a homogeneous-scale transition model could compensate for this $a_{B}$ discrepancy within the DESY5 compilation, the same model would, in turn, jeopardize the original $a_{B}$ consistency of PantheonPlus and reproduce a new $a_{B}$ discrepancy within the PantheonPlus compilation. In addition, the 1D probability density distribution (red dashed line) of low-$z$ intercept $-5a_{B}$ in Fig. 2 is non-Gaussian, unlike the other two cases, suggesting the low-$z$ SNe suffer from large scatters and do not follow a consistent magnitude-distance relation. As a comparison, the PantheonPlus SNe contain even more subsamples (18 samples) but still converge to a Gaussian distribution around the weighted average, indicating their well-corrected SNe have followed an accordant magnitude-distance relation just as the DES-SN does. Therefore, the strange transition and large scatter in the low-$z$ intercept $-5a_{B}$ can only be attributed to systematics (including an $M_{B}$ transition). Regardless of its origin, these low-$z$ SNe can only be used for cosmological analyses after correcting this mismatch intercept.
3.3 Systematics corrections
The low-$z$ SNe systematics call for better sample selections and calibrations, which goes well beyond our current study. Based on the $a_{B}$ diagnosis, we propose a simple but rough correction procedure to the low-$z$ sample, which is, however, sufficient for our purpose to re-examine the preference for dynamic DE. Since the intercept $-5a_{B}$ measured from high-$z$ SNe is less affected by systematics from the peculiar velocity effect than that measured from low-$z$ SNe, the systematics correction for low-$z$ SNe is done by matching the low-$z$ intercept to the high-$z$ intercept. Specifically, our correction procedures include three kinds: (i) Smoothing the irregular fluctuations in the intercept $-5a_{B}$ for each redshift bin of low-$z$ SNe to match the weighted average of all low-$z$ SNe,
$m_{B,i\in\mathrm{low}-z}^{\mathrm{fluc-debias}}:=5\lg\hat{d}_{L}(z_{i\in\mathrm{low}-z})-5\overline{a_{B,\mathrm{low}-z}}\,;$ (3)
(ii) Reducing the low-$z$ $m_{B}$ by $0.043$ magnitude so that the weighted average $-5a_{B}$ of low-$z$ SNe also matches the weighted average of DES-SN,
$m_{B,i\in\mathrm{low}-z}^{\mathrm{disp-debias}}$ $:=5\lg\hat{d}_{L}(z_{i\in\mathrm{low}-z})-5a_{B,i\in\mathrm{low}-z}-0.043$
$\equiv m_{B,i\in\mathrm{low}-z}^{\mathrm{std}}-0.043\,;$ (4)
(iii) Combining correction processes (i) and (ii) together to have a fully de-biased low-$z$ SN sample,
$m_{B,i\in\mathrm{low}-z}^{\mathrm{fully-debias}}$ $:=5\lg\hat{d}_{L}(z_{i\in\mathrm{low}-z})-5\overline{a_{B,\mathrm{low}-z}}-0.043$
$\equiv 5\lg\hat{d}_{L}(z_{i\in\mathrm{low}-z})-5\overline{a_{B,\mathrm{DES-SN}}}.$ (5)
As a preliminary test, we find a $\sim 2.2\sigma$ preference for dynamical DE from CMB+original low-$z$ SNe decreases to $\sim 1.5\sigma$ after roughly de-biasing the fluctuations in $-5a_{B}$ as shown in the first group of constraints in Tab. 2, which also slightly regularizes the $H_{0}-M_{B}$ contour (black unfilled contours) in the inset of the right panel of Fig. 1. We further test our correction procedures when including DESI Y1/Y3 BAO as shown in the upper/lower-left panels of Fig. 3, respectively, as well as the second/fourth groups of constraints in Tab. 2. In the upper/lower-left panel of Fig. 3 with inclusions of DESI Y1/Y3 BAO, correcting for the fluctuations in $-5a_{B}$ alone can cause little deviation in the $w_{0}-w_{a}$ contour, however, solely correcting for the low-$z$ discrepancy in $-5_{B}$ without debiasing the fluctuations first is already sufficient to reduce the preference for dynamical DE from $\sim 3.5\sigma/3.7\sigma$ to $\sim 1.6\sigma$, almost overlapping with the fully de-biased correction case with $\sim 1.5\sigma/1.6\sigma$. A larger reduction to $\sim 0.5\sigma/0.7\sigma$ and the same pattern also emerges in the upper/lower-right panels of Fig. 3 with DESI Y1/Y3 BAO when further excluding Planck-CMB assuming fiducial $H_{0}=67.4$ km/s/Mpc as shown in the third and fifth groups of constraints in Tab. 2. A different $H_{0}$ prior only changes the degenerated constraints with respect to the sound horizon $r_{d}$ and absolute magnitude $M_{B}$ but still preserves the same reduced tension in the $w_{0}-w_{a}$ plane for this purely late-Universe data combination (DESI BAO+DESY5) without early-Universe Planck-CMB calibration. Therefore, it is the $\sim 0.043$ mag. discrepancy in $-5a_{B}$ between low-$z$ and DES-SN that plays a leading role in driving the preference for dynamical DE.
4 Conclusions and Discussions
Recent cosmological analyses from DESI DR1 and DR2 have attracted wide interest but also raised significant challenges and hot debate to the standard $\Lambda$CDM. In this letter, we reduce the challenge by locating some systematics at low-redshift supernovae. The multi-sample mixed low-$z$ SN sample in DESY5 compilation, suffering from large scatters, fails to rebuild a consistent intercept of the magnitude-distance relation and admits a $\sim 0.043$ magnitude discrepancy with respect to the DES-SN at higher redshifts, which largely drives the preference for dynamical DE. After the first and second versions of our preprint, the DESI DR2 paper (M. Abdul Karim et al., 2025a) has explicitly tested this observation by directly getting rid of low-$z$ SNe in CMB+DESI DR2+DESY5($z>0.1$), where the preference for dynamical DE is reduced to be $2\sigma$ as shown in the middle panel of Fig. 14. After roughly correcting this discrepancy in the intercept for the low-$z$ SN samples, we find the corrected datasets (Planck+DESI DR1/DR2+DESY5) favor dynamical DE only at $1.5\sigma/1.6\sigma$ levels, not significant at all to call for dynamical DE. Therefore, the success of the standard $\Lambda$CDM has not been overturned.
Although we have identified the systematics involving $a_{B}$ inconsistency in the low-$z$ SN sample within the single DESY5 compilation is largely responsible for the immature claim of dynamical DE, the DESI DR1 BAO data alone indeed shows some deviations from the Planck-$\Lambda$CDM model at higher redshifts, especially for those luminous red galaxies (LRG) between $0.80.6<z<$ 0.8 (A. G. Adame et al., 2025a), which have also been considered recently by Z. Wang et al. (2024); G. Liu et al. (2024) as the possible drives for the preference of dynamical DE. This SN-independent deviation from Planck-$\Lambda$CDM model has been re-enforced recently in the DESI DR2 papers (M. Abdul Karim et al., 2025a, b; K. Lodha et al., 2025; W. Elbers et al., 2025) with an explicit discrepancy in $\Omega_{m}$ between Planck-CMB and DESI-BAO and a clear deviation from Planck-$\Lambda$CDM model for LRG2 BAO measurements. Whether this DESI-Planck tension should be better explained by dynamical DE (with phantom crossing) is optional (Z.-K. Guo et al., 2005; B. Feng et al., 2005, 2006), as it can be equally accounted for by some modified gravity with non-minimal coupling (G. Ye et al., 2024; J. Pan & G. Ye, 2025; W. J. Wolf et al., 2025b), mimicking thawing DE (S.-F. Chen et al., 2024; W. J. Wolf et al., 2025a, 2024) even with phantom crossing (R.-G. Cai et al., 2021; A. Chakraborty et al., 2025; J. Khoury et al., 2025).
We thank Tamara Davis, George Efstathiou, Bin Hu, Yipeng Jing, Cheng Zhao, Gong-Bo Zhao, Jian-Qi Liu, and Yan-Hong Yao for helpful discussions. This work is supported by the National Key Research and Development Program of China Grants No. 2021YFA0718304, No. 2021YFC2203004, and No. 2020YFC2201501, the National Natural Science Foundation of China Grants No. 12422502, No. 12105344, No. 11821505, No. 12235019, No. 11991052, No. 12447101, No. 12073088, and No. 11947302, the China Manned Space Program with Grant No. CMS-CSST-2025-A01, and the Postdoctoral Fellowship Program of CPSF. We also acknowledge the use of the HPC Cluster of ITP-CAS. {contribution} All authors contributed equally to the manuscript.